Diffusion Morning Report 10/18/04

tberg tberg at snf.stanford.edu
Mon Oct 18 07:37:32 PDT 2004


Good Morning All,
    Looking pretty good this morning.  Nitride and poly are under 
observation.  LTOs pump was off-looking into it.  Tylan 2 had temp
issues -remade TC connections-ok now.  WBsilicide shutdown for decon in 
prep for move.  Epi regulator is being rebuilt-hope for return this week.
Everything else looking OK for now.

> . 	Date: 	10/18/2004 	 
> Tube Name/Number 	Status 	Comments 	THICKNESS-Change at
> Tylan 1 	Up 	  	 
> Tylan 2 	Up 	  	 
> Tylan 3 	Up 	  	 
> Tylan 4 	Up 	  	 
> Tylan 5-BBr3 	Up 	  	 
> Tylan 6-POCl3 	Up 	  	 
> Tylan 7-TEOS2 	down 	Qual testing 	 
> Tylan 8-SiGe 	Up 	  	12.33/30
> Tylan 9-Poly 	yellow 	observation 	11.87//30
> Tylan 10-Nitride 	yellow 	possible particle issues 	2.79/30
> Tylan 11-BPSG-LTO 	down 	Pump was off investigating 	21.1/30
> Tylan 12-FGA 	Up 	  	 
> Tystar 	Up 	  	7.46/30
> Doped Poly 	GONE 	  	 
> TEOS 	up 	  	 
> Tylan 16 	N/A 	  	 
> Thermco 1 	Up 	  	 
> Thermco 2 	Up 	  	 
> Thermco Poly 	yellow 	Testing 	 
> Thermco Nitride 	Down 	Waiting 	 
>   	  	  	 
> Burn Box 	Up 	  	 
> ZYGO 	up 	  	 
>   	  	  	 
> GAAS 22 	UP 	  	 
> GAAS 23 	UP 	  	 
>   	  	  	 
> SiH4 for furnaces 	  	Auto changeover working well 	B on line 520 PSI
>

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20041018/1c57bc55/attachment.html>


More information about the diffusion mailing list